OpenStudy (anonymous.):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (texaschic101):

do you know the slope of y = 1/2x - 8 ?

OpenStudy (texaschic101):

In y = mx + b form (which is what the equation is in) the m position is the slope. So what is the slope ?

OpenStudy (anonymous.):

Yea ! y-6 =1/2(x-2)

OpenStudy (texaschic101):

almost..... y - 6 = 1/2(x -(-2) y - 6 = 1/2(x + 2) y - 6 = 1/2x + 1 y = 1/2x + 1 + 6 y = 1/2x + 7 any questions ?

OpenStudy (anonymous.):

No :D !! Thank you soo much for your help :)

OpenStudy (texaschic101):

anytime :)